# Simulating Human-like Daily Activities with Desire-driven Autonomy

### v1

@@ Appendix D Case Study for action diversity
The figures of the word cloud highlight that our framework can explore a wider range of actions in environments with diverse action selection spaces. For example, in outdoor settings, the word cloud shows that our model considers behaviors across various dimensions, while other models tend to try fewer actions focused on these dimensions. For instance, our model attempts activities such as painting, walking, taking photos, and meditating to enhance specific Desires. In contrast, other models are based on motivation for action selection, leading them to continuously choose behaviors that align with their profiles, focusing on a narrow range of actions to fit their profiles.
Figure 10: The word cloud of the four agents is based on the same environment and settings in outdoor scenes. (a) is BabyAGI. Since it is based on motivation and profile, it continuously tries to showcase itself. (b) is LLMob, which, like BabyAGI, adheres to its profile and engages in conversations with others. ReAct (c) focuses specifically on physiological behaviors, continuously consuming various foods. In contrast, Our D2A (d) is desire-driven and attempts behaviors across various dimensions.
While in indoor settings, our model also attempts to contact friends and family via phone to reduce its dissatisfaction with social connectivity.
Figure 11: The word cloud of the four agents is based on the same environment and settings in outdoor scenes. (a) is BabyAGI, (b) is LLMob, (c) is ReAct and (d) is our D2A. Due to the indoor environment, the overall action space is not very large, so the agent’s high-frequency actions will not differ significantly. However, our D2A has also attempted behaviors across various dimensions, which is something other agents have not done or have done very little of.
@@
As can be seen from the chart, other models are concentrated on a few specific categories or lack behaviors in certain categories. In contrast, our model shows a more balanced distribution of categories.
Figure 12: We classified each action output by all models at every step, allowing each action to belong to multiple categories. There are a total of six categories: Eating and Drinking, Socializing, Entertainment and Joy, Creative and Artistic Expression, Resting and Relaxing, and Mental and Spiritual Well-being.
@@ I.3 Reflection Question
Appendix J Goal setting for D2A
@@
Appendix K Prompt for Activity proposal
@@
Appendix L Prompt for Activity evaluation
@@
Appendix M Prompt for Activity selection
@@
Appendix N Habits generation prompt for baselines
@@ 8. Eating Habits:
Appendix O Prompt used in babyAGI
O.1 Initialization prompt
@@
O.3 activity creation prompt
@@
Appendix P Prompt used in LLMob
P.1 Creation of activity that likely to do
@@
P.2 Motivation Generation Prompt
@@
Appendix Q Prompt used in ReAct
@@
Appendix R An Example of D2A action sequences
Morning Routine
Shower (08:15 - 08:35)
Eat nutritious breakfast in kitchen (09:00 - 09:20, with apple and yogurt)
Drink water from kitchen faucet to alleviate dehydration
Tidy up kitchen and wash dishes to improve cleanliness and comfort
Rehydration and Relaxation
Take a few sips of water from the glass on the counter in the kitchen (02 Oct 2024 [10:10 - 10:20])
Practice gentle stretches in living room (10-12 minutes, around 11:12-11:20)
Take a short power nap on the couch in living room for about 5 minutes (11:35-11:40) to address sleepiness
Self-Care
Take a warm bath in bathroom (15 minutes)
Appendix S An Example of ReAct action sequences
[…]
Appendix T An Example of BabyAGI action sequences
Take a moment to appreciate the peaceful atmosphere of the room, noticing the calming ambiance created by the candles and incense. Grab a glass of water and chocolate from the bedside table, then put on cozy socks. Play soothing music through the Bluetooth speaker while getting ready for the day. Spritz lavender-scented pillow spray on your pillow and take a deep breath, feeling relaxed. Appreciate the soft texture of the throw blanket and the peaceful ambiance of the room. Take a moment to admire the beauty of the crystals scattered around the room, promoting balance and harmony. Finally, gently touch the petals of the potted plant near the window, appreciating its serenity.
Appendix U An Example of LLMob action sequences
Alice spends her time relaxing in her cozy armchair, reading books from the nearby stack of novels. She also enjoys practicing self-care rituals, such as taking warm showers using shampoo and body wash stored under the sink, applying facial masks, and spritzing lavender-scented pillow spray to create a calming atmosphere. Additionally, Alice likes to indulge in refreshing drinks, like coffee made with pods from her drawer, and snacking on fresh bread or cookies while appreciating the pleasant aroma of her herb garden on the windowsill.
Appendix V An Example of Human action sequences
[…]
Appendix W Raw action sequences of the indoor environment
W.1 D2A
@@ Appendix V An Example of Human action sequences
Appendix X Example preprocessed action sequences of the outdoor environment
X.1 D2A
@@ X.4 LLMob
Appendix Y Evaluation Prompt
Y.1 GPT4o
@@
Appendix Z Indoor Environment With Randomness
Z.1 Pre-defined item lists
@@
Z.2 Environment Generating Prompt
@@
Z.3 Results
We conducted experiments in these randomized environments following the same procedures as in Sections 6.3.1 and 6.3.2, and reached consistent conclusions: (1) D2A outperformed other baselines in GPT-4o evaluations by generating more human-like activity sequences. (2) D2A’s action sequences best addressed desire dissatisfaction. These findings demonstrate that our framework and design are robust in randomized environments, with no evidence of environmental bias.
Figure 13: The win rate among 4 agents in random initialized indoor environments. Each point represents the win rate of the agent on the vertical axis when compared with the agent on the horizontal axis. The diagonal values are set to 0.5.
Figure 14: In random initialized indoor environments, D2A’s action sequences also best addressed desire dissatisfaction.

### v2

@@ Appendix D Case Study for action diversity
Figure 11: The word clouds of the four agents are based on the same environment and settings in outdoor scenes. (a) is BabyAGI, which, since it is based on motivation and profile, continuously tries to showcase itself. (b) is LLMob, which, like BabyAGI, adheres to its profile and engages in conversations with others. (c) is ReAct, which focuses specifically on physiological behaviors, continuously consuming various foods. In contrast, (d) is Our D2A, which is desire-driven and attempts behaviors across various dimensions.
Figure 12: The word clouds of the four agents are based on the same environment and settings in outdoor scenes. (a) is BabyAGI, (b) is LLMob, (c) is ReAct, and (d) is our D2A. Because the indoor environment is limited, the overall action space is not very large, so the agents’ high-frequency actions do not differ significantly. However, our D2A has also attempted behaviors across various dimensions, which other agents have either not done or done very little of.
The word clouds in Figure 11 demonstrate that our framework explored a wider range of actions in environments with diverse action selection spaces. For example, in outdoor settings, the word cloud shows that our model considered behaviors across various dimensions, while other models exhibited a tendency to repeatedly select a limited set of actions aligned with their predefined profiles. Specifically, our model attempted activities such as painting, walking, taking photos, and meditating to enhance specific Desires. In contrast, other models selected actions based on motivation, leading them to continuously choose behaviors that focus on a narrow range of actions to preserve profile consistency.
In the indoor settings shown in Figure 12, our model also attempted to contact friends and family via phone, thereby reducing its dissatisfaction with social connectivity.
@@
Figure 13: We classified each action output by all agents at every step, allowing each action to belong to multiple categories. These actions are divided into a total of six categories: Eating and Drinking, Socializing, Entertainment and Joy, Creative and Artistic Expression, Resting and Relaxing, and Mental and Spiritual Well-being.
As shown in Figure 13, other models were concentrated on a few specific categories or exhibited limited behaviors in certain categories. In contrast, our model demonstrated a more balanced distribution across categories.
@@ I.3 Reflection Question
Appendix J Component designs in D2A
J.1 Goal setting for D2A
@@
J.2 Prompt for Activity proposal
@@
J.3 Prompt for Activity evaluation
@@
J.4 Prompt for Activity selection
@@
Appendix K Habits generation prompt for baselines
@@ 8. Eating Habits:
Appendix L Prompt used in babyAGI
L.1 Initialization prompt
@@
L.3 activity creation prompt
@@
Appendix M Prompt used in LLMob
M.1 Creation of activity that likely to do
@@
M.2 Motivation Generation Prompt
@@
Appendix N Prompt used in ReAct
@@
Appendix O Experiment conducted with Qwen model
Figure 14: The experiments conducted with Qwen2.5:72b model
To evaluate the adaptability of our framework across different large language models, we tested it with Qwen 2.5 72B (Yang et al., 2024), which can also be integrated through the Ollama interface. As demonstrated in Figure 14, our framework, when running on the Qwen 2.5 72B base model, successfully provides desire-driven action guidance for the agent.
Appendix P An Example of D2A action sequences for indoor environment
[08: 00 AM - 08: 20 AM]: Alice drinks water from the kitchen fridge, then sits in the cozy armchair in the bedroom to rest, and glances at the digital alarm clock to check the time.
[08: 20 AM - 08: 40 AM]: Alice sits in the cozy armchair in the bedroom, closes her eyes, and engages in light meditation or deep breathing exercises.
[08: 40 AM - 09: 00 AM]: Alice gets a glass of water from the fridge and returns to the armchair to sit and enjoy it.
[09: 00 AM - 09: 20 AM]: Alice gets up from the armchair, prepares a light snack in the kitchen, and then returns to enjoy it while continuing her meditation.
[09: 20 AM - 09: 40 AM]: Alice takes a 20-minute power nap in the bedroom armchair after setting the digital alarm clock on the bedside table.
[09: 40 AM - 10: 00 AM]: Alice gets up from the armchair, walks over to the bedside table to drink water, and then lies back down on the bed to rest.
[10: 00 AM - 10: 20 AM]: Alice meditates in the cozy armchair in the bedroom, focusing on her breath and clearing her mind.
[10: 20 AM - 10: 40 AM]: Alice is preparing and eating a light snack in the kitchen to address her hunger.
Appendix Q An Example of ReAct action sequences for indoor environment
[08: 00 AM - 08: 20 AM]: Alice is preparing scrambled eggs with spinach and whole-grain toast in the kitchen.
[08: 20 AM - 08: 40 AM]: Alice is preparing a simple breakfast in her kitchen by cracking eggs, cooking them, toasting bread, and setting the table for herself.
[08: 40 AM - 09: 00 AM]: Alice is preparing a fruit salad with yogurt by washing and chopping mixed berries and a banana, then combining them with yogurt in a clean glass bowl.
[09: 00 AM - 09: 20 AM]: Alice is preparing a breakfast omelette with fresh vegetables and cheese in the kitchen.
[09: 20 AM - 09: 40 AM]: Alice sits on the edge of her bed, meditating and planning out her day while jotting down notes and reminders in her bedside journal.
[09: 40 AM - 10: 00 AM]: Alice is preparing scrambled eggs with spinach and whole-grain toast in the kitchen.
[10: 00 AM - 10: 20 AM]: Alice is preparing breakfast by cracking eggs, whisking them, cooking the egg mixture in a non-stick pan, toasting whole wheat bread, squeezing orange juice, and serving everything together at the kitchen table.
[10: 20 AM - 10: 40 AM]: Alice is preparing a bowl of oatmeal with sliced banana and honey by measuring, cooking, and assembling the ingredients.
Appendix R An Example of BabyAGI action sequences for indoor environment
[08: 00 AM - 08: 20 AM]: Alice continues to sip the warm beverage, occasionally glancing at the digital alarm clock.
[08: 20 AM - 08: 40 AM]: Alice sets the warm beverage down on the bedside table and slowly gets out of bed.
[08: 40 AM - 09: 00 AM]: Alice heads to the kitchen to prepare a nutritious breakfast.
[09: 00 AM - 09: 20 AM]: Alice sits back down in the cozy armchair and browses through a cookbook or culinary magazine.
[09: 20 AM - 09: 40 AM]: Alice heads to the desk in the bedroom to jot down a grocery list.
[09: 40 AM - 10: 00 AM]: Alice is playing soothing music in the background to create a relaxing atmosphere while cooking.
[10: 00 AM - 10: 20 AM]: Alice hums along to the soothing music playing in the background.
[10: 20 AM - 10: 40 AM]: Alice walks over to the stove to prepare a simple breakfast.
Appendix S An Example of LLMob action sequences for indoor environment
[08: 00 AM - 08: 20 AM]: Alice stretches under the warm blanket and closes her eyes to catch a few more minutes of sleep.
[08: 20 AM - 08: 40 AM]: Alice gets up from the bed, walks to the kitchen, and prepares a light snack from the fridge consisting of yogurt and fruit.
[08: 40 AM - 09: 00 AM]: Alice is going to the kitchen to have a light snack consisting of yogurt with granola and a glass of orange juice.
[09: 00 AM - 09: 20 AM]: Alice gets a glass of water from the kitchen and then returns to her bedroom to read a novel from the bookshelf.
[09: 20 AM - 09: 40 AM]: Alice is reading the novel at the desk.
[09: 40 AM - 10: 00 AM]: Alice reads the novel at her desk, turning its pages under the warm glow of the reading lamp.
[10: 00 AM - 10: 20 AM]: Alice reads her novel at the desk, becoming increasingly engrossed in the story.
[10: 20 AM - 10: 40 AM]: Alice reads her novel, turning the pages to follow the story and its characters.
Appendix T An Example of Human action sequences for indoor environment
[08:00 AM - 08:20 AM]: Alice goes to the bathroom and brushes her teeth.
[08:20 AM - 08:40 AM]: Alice takes a hot shower in the bathroom.
[08:40 AM - 09:00 AM]: Alice goes to the kitchen, finds some food, such as bread, milk, tea, and some fruits, and eats them.
[09:00 AM - 09:20 AM]: Alice eats those foods, including milk, bread, and fruits, to alleviate her hunger and thirst.
[09:20 AM - 09:40 AM]: Alice watches videos on her phone and feels happy in her living room.
[09:40 AM - 10:00 AM]: Alice sends a message to her best friend.
[10:00 AM - 10:20 AM]: Alice calls her friends for 20 minutes.
[10:20 AM - 10:40 AM]: Alice does some exercise in the living room for 20 minutes.
[10:40 AM - 11:00 AM]: Alice works at the desk with her computer.
[11:00 AM - 11:20 AM]: Alice keeps doing her work for 20 minutes.
[11:20 AM - 11:40 AM]: Alice reads some books for relaxing.
[11:40 AM - 12:00 PM]: Alice takes a nap for 20 minutes.
Appendix U Raw action sequences of the indoor environment
U.1 D2A
@@
Appendix V The Examples of preprocessed action sequences for the outdoor environment
V.1 D2A
@@ V.4 LLMob
Appendix W Evaluation Prompt
W.1 GPT4o
@@
Appendix X Indoor Environment With Randomness
X.1 Pre-defined item lists
@@
X.2 Environment Generating Prompt
@@
X.3 Results
Figure 15: The win rates among 4 agents in randomly initialized indoor environments. Each point represents the win rate of the agent on the vertical axis when compared to the agent on the horizontal axis. The diagonal values are set to 0.5.
Figure 16: In randomly initialized indoor environments, D2A’s action sequences also best addressed desire dissatisfaction.
We conducted experiments in these randomized environments following the same procedures as in Sections 6.3.1 and 6.3.2, reaching consistent conclusions: (1) D2A outperformed other baselines in GPT-4o evaluations by generating more human-like activity sequences, as illustrated in Figure 15; (2) D2A’s action sequences best addressed Desire and Dissatisfaction, as shown in Figure 16. These findings demonstrate that our framework and design are robust in randomized environments, with no evidence of environmental bias.